# Finiteness of totally geodesic hypersurfaces

### v2

@@ 3.1.1. Diagram of vectors and subspaces
Let $UM\xrightarrow{\pi_{1}}M$ denote the unit sphere bundle, $HM\xrightarrow{\pi_{n-1}}M$ the Grassmannian bundle of oriented hyperplanes, i.e. $(n-1)$-dimensional tangent planes, and $FM$ the bundle of pairs $(v,\Pi)$ where $v$ is unit vector, $\Pi$ is an oriented hyperplane, and $v\in\Pi$. Denote the forgetful projections by $FM\xrightarrow{\pi_{U}}UM$ and $FM\xrightarrow{\pi_{H}}HM$.
@@
(3.1.1) ${FM}$${UM}$${HM}$${M}$$\scriptstyle{\pi_{U}}$$\scriptstyle{\pi_{H}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{n-1}}$      ${I}$${A}$${Z}$$\scriptstyle{\pi_{U}}$$\scriptstyle{\pi_{H}}$
Our notation is summarized in (missing) 3.1.1. Typically, a point in $M$ will be denoted $p$, a point in $HM$ will be denoted $(p,\Pi)$ with $\Pi\subset T_{p}M$, and a point in $UM$ will be denoted $(p,v)$ with $v\in T_{p}M$.
@@ 3.2.1. Setup
We keep the notation of § 3.1.2. It is immediate from the definitions that $Z\subset HM$ is a closed subset, since a totally geodesic hypersurface $S\ni p$ with $T_{p}S=\Pi$ must locally be equal to the exponential map of a ball in $\Pi$, and a hypersurface is totally geodesic if and only if its second fundamental form vanishes identically (see [O’N83, §12-13]).
We will call a set real-analytic if it is locally the vanishing locus of finitely many real-analytic functions. A subset is subanalytic if it is locally the image, under a real-analytic map, of a relatively compact real-analytic subset (see [BM88, Def. 3.1]). Many equivalent descriptions of subanalytic sets exist, see [BM88, Prop. 3.13].
@@
The statement is local (by connectedness of the fibers) so by an appropriate choice of coordinates on $X$ we can assume that $\pi$ is a coordinate projection in a ball of $\mathbb{R}^{m+n}$ to $\mathbb{R}^{m}$. Using a multi-index notation with $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, we can expand in Taylor series: $\rho(\mathbf{x},\mathbf{y})=\sum_{I}a_{I}(\mathbf{x})\mathbf{y}^{I}$ and then $Z$ is given as the vanishing locus of all the $a_{I}(\mathbf{x})$. However, by [Nar66, Ch. V, Cor. 2, pg. 100] it follows that $Z$ can be given as the vanishing locus of finitely many real-analytic functions in $\mathbf{x}$. ∎
@@ 3.2.4. Structure of real-analytic sets
By [BM88, 2.11-12] an analytic set $Z$ has a stratification into finitely many locally closed pieces that are analytic submanifolds. The maximal dimension of a stratum is called the dimension of $Z$.
@@
Let $A^{\circ}\subset A$ denote the set of smooth points, i.e. it is real-analytic, with $\dim A^{\circ}=\dim A$, and $\dim\left(A\setminus A^{\circ}\right)<\dim A$ see e.g. [BM88, Theorem 7.2]. Consider the map $\pi_{U}\colon I\to A$. Either $\dim A=\dim I$ (call this Case 1), or there exists a further relatively open subset $A^{\circ\circ}\subset A^{\circ}$ such that the map $\pi_{S}|_{I}$ has fibers of positive-dimension over $A^{\circ\circ}$ (call this Case 2).
@@
We have that $T_{a}UM=W^{s}(a)\oplus W^{c}(a)\oplus W^{u}(a)$ and $W^{c}(a)\subset T_{a}A$, since it represents the geodesic flow direction. In the quotient $T_{a}UM/W^{c}(a)$ the hyperplane $T_{a}A/W^{c}(a)$ must be transverse to at least one of the images of $W^{c/s}(a)$, say $W^{s}(a)$. Therefore the intersection $W^{cs}(a)\cap T_{a}A$ must have the expected dimension, so we have the equalities
@@
Pick an $(n-1)$-dimensional germ of smooth real-analytic hypersurface $N\subset A^{\circ}$ with $a\in N$ such that $T_{a}UM=T_{a}N\oplus W^{cs}(a)$. For example, $N$ can be obtained by intersecting $A^{\circ}$ with a linear space of the appropriate dimension in local coordinates. We thus have that in a neighborhood of $a$, the manifold $N$ is transverse to the center-stable foliation.
@@
The map $\pi_{1}\colon\mathcal{W}^{cs}_{loc}[a]\to M$ is a diffeomorphism onto its image (see for example [BP07, §12.2]). Therefore, there exists an inverse map $\sigma_{a}\colon U\to\mathcal{W}^{cs}_{loc}[a]$ which is again a diffeomorphism onto its image. For $p\in U$, the unit tangent vector $\sigma_{a}(p)$ is characterized by the property that it is the unique vector above $p$ such that $g_{t}(\sigma_{a}(p))$ and $g_{t}a$ remain at bounded distance as $t\to+\infty$.
@@ 3.3. Analyzing the transitivity group
As we saw in § 2.2.14, there exists a continuous reduction of the structure group of the frame bundle to the transitivity group. Restricting this reduction to the unit sphere over one point of $M$ already yields nontrivial obstructions, first exploited by Brin and Gromov [BG80]. A systematic analysis of known obstructions is given in [CLMS23, §3.3], which we use to extract the needed constraints on the transitivity group. While ergodicity of the frame flow is not known in the generality we would need, the existing restrictions coupled with additional arguments are sufficient for our purposes.
@@ 3.3.1. The simplest obstruction
Recall that even-dimensional spheres have Euler characteristic $2$, hence do not admit a nowhere vanishing vector field, or even an everywhere defined tangent line-field. This is originally due to Poincaré in dimension $2$ [Poi85, Chapter 13] and Brouwer in higher dimensions [Bro12]. In other words, if $n\geq 3$ is odd, there does not exist a continuous map
@@ 3.3.2. Surjectivity of image
We also record the following elementary consequence for even $n$. Suppose $\beta\colon\mathbb{S}^{n-1}\to\mathbb{S}^{n-1}$ is continuous, with the property that $\beta(e)\perp e$. Then $\beta$ is surjective when its image is descended to $\mathbb{R}\mathbb{P}^{n-1}:=\mathbb{S}^{n-1}/\pm 1$. Indeed, suppose by contradiction that the two vectors ±$\pm e_{n}$ are never in the image of $\beta$. Restricted to $\mathbb{S}^{n-2}(e_{n}^{\perp})$ we can decompose $\beta(e)=\beta_{1}(e)+\beta_{2}(e)$ with $\beta_{2}(e)=\alpha(e)e_{n}$ and $\beta_{1}(e)\in e_{n}^{\perp}$. Now by assumption $\beta_{1}(e)\neq 0$ and also $\left\langle e,\beta_{1}(e)\right\rangle=0$, so we can define
@@
which has the property that $\widetilde{\beta}(e)\perp e$. But this is a contradiction to § 3.3.1 since $n-2$ is even.
@@ 3.3.3. Restrictions on reduction of structure group
We next recall some more elaborate topological facts from [CLMS23, Proof of Thm. 3.1, Step 1]. Consider the principal $\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$-bundle over the $(n-1)$-sphere:
@@ 3.3.4. Full frame bundle
We extend (missing) 3.1.1 to include the full frame bundle:
(3.3.2) ${PM}$${FM}$${UM}$${HM}$${M}$$\scriptstyle{\pi_{U}}$$\scriptstyle{\pi_{H}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{n-1}}$
@@
(3.3.3) $\begin{split}PM(x)&:=\left\{(e_{1},\dots,e_{n})\hookrightarrow T_{x}M\text{ an oriented frame}\right\}\\ UM(x)&:=\left\{e_{1}\hookrightarrow T_{x}M\text{ unit vector}\right\}\\ HM(x)&:=\left\{e_{n}\hookrightarrow T_{x}M\text{ unit vector}\right\}\\ FM(x)&:=\left\{(e_{1},e_{n})\hookrightarrow T_{x}M\text{ a pair of unit vectors with }e_{1}\perp e_{n}\right\}\end{split}$
@@
With assumptions as in Theorem 1, let $H\subset\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$ be the transitivity group of $PM\to UM$, which we assume is connected by § 2.2.14. Then we have the following possibilities for $H$ and coset space $C:=\raisebox{-1.00006pt}{$H$}\backslash\raisebox{1.00006pt}{$\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$}/\raisebox{-1.00006pt}{$\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$}$:
@@
    $H=\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ and $C=[-1,1]$ (any $n$).
@@
    $H=\operatorname{\mathbf{SU}}(3)$ and $C=[-1,1]$ (for $n=8$ only).
Above, “any $n$” means subject to $n\geq 3$ and the specified parity condition.
@@
Suppose next that $n$ is even. Because a hypersurface $S$ is necessarily odd-dimensional, we have that $H_{S}=\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ except possibly when $n=8$ in which case $H_{S}=\operatorname{\mathbf{SU}}(3)$ is also possible. So for all even $n\neq 8$ we have the claim, and since the orbits of $\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ on $\mathbb{S}^{n-2}\subset\mathbb{R}^{n-1}$ are classified by their last coordinate, an element of $[-1,1]$. For $n=8$ we also have the possibility that $H=H_{S}=\operatorname{\mathbf{SU}}(3)$. But the corresponding action of $\operatorname{\mathbf{SU}}(3)$ on $\mathbb{S}^{6}$ has orbit space again equal to $[-1,1]$ and given by the value of the last coordinate. ∎
@@
From (missing) 3.1.1 it suffices to show that the set $I(x)\subset FM(x)$ has the same dimension as $FM(x)$, since $I=\pi_{H}^{-1}(Z)$. Recall also that $FM=PM/\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ and we have an action of $G_{t}$ lifting the geodesic flow $g_{t}$ on $UM$. Let $H\subset\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$ be the transitivity group of $PM\to UM$.
@@
[…]
@@
From § 3.3.4, only the cases when $C\neq\left\{pt\right\}$ need to be treated, i.e. $C=[-1,1]$ and $n$ is even.
@@
is a nonempty open set, as the intersection of two open spherical caps around $\beta(e_{1}^{\prime})$ and $-\beta(e_{1}^{\prime\prime})$. Non-emptiness holds because the two spherical caps have measure adding up to the measure of $\mathbb{S}^{n-1}$, and the only situation in which they are disjoint is when $\beta(e_{1}^{\prime})=\beta(e_{1}^{\prime\prime})$. For a continuous path $e_{1}(t)$ with $e_{1}(0)=e_{1}^{\prime},e_{1}(1)=e_{1}^{\prime\prime}$, we have that any element $e_{n}\in L$ is in some $\mathbb{S}^{n-2}_{\alpha}(e_{1}(t))$, since $f(t):=\left\langle\beta(e_{1}(t),e_{n})\right\rangle$ satisfies $f(0)>\alpha,f(1)<\alpha$, so by the intermediate value theorem there exists $t$ such that $f(t)=\alpha$. This shows that an open set of hyperplanes belongs to $Z$ and concludes the proof. ∎
@@ 3.3.7. Last step
We know that $Z\subset HM$ is a closed real-analytic set of full dimension, therefore it is in fact all of $HM$. Indeed, for any point of $z\in Z$, the germ of $Z$ at $z$ denoted $Z_{z}$ has full dimension, therefore by [Nar66, Ch. V, Prop. 1, pg. 91] the complexified germ $\widetilde{Z}_{z}$ is an entire complex neighborhood of $z$, and hence the germ $Z_{z}$ is an entire real neighborhood of $z$ by the result just cited. It follows that $Z$ is also open, hence equal to $HM$.
It now follows from “Cartan’s axioms of $k$-planes” (with $k=n-1$) Theorem 2, that $M$ has constant sectional curvature. Since the geodesic flow is Anosov, the constant curvature must be negative.

### v3

@@ 3.1.1. Diagram of vectors and subspaces
Let $UM\xrightarrow{\pi_{1}}M$ denote the unit sphere bundle, $HM\xrightarrow{\pi_{n-1}}M$ the Grassmannian bundle of oriented hyperplanes, i.e. (n1)(n-1)-dimensional tangent planes, and $FM$ the bundle of pairs $(v,\Pi)$ where $v$ is unit vector, $\Pi$ is an oriented hyperplane, and $v\in\Pi$. Denote the forgetful projections by $FM\xrightarrow{\pi_{U}}UM$ and $FM\xrightarrow{\pi_{H}}HM$.
@@
[…]
Our notation is summarized in () 3.1.1. Typically, a point in $M$ will be denoted $p$, a point in $HM$ will be denoted $(p,\Pi)$ with $\Pi\subset T_{p}M$, and a point in $UM$ will be denoted $(p,v)$ with $v\in T_{p}M$.
@@ 3.2.1. Setup
We keep the notation of § 3.1.2. It is immediate from the definitions that $Z\subset HM$ is a closed subset, since a totally geodesic hypersurface $S\ni p$ with $T_{p}S=\Pi$ must locally be equal to the exponential map of a ball in $\Pi$, and a hypersurface is totally geodesic if and only if its second fundamental form vanishes identically (see [ONeill1983_Semi-Riemannian-geometry, §12-13]).
We will call a set real-analytic if it is locally the vanishing locus of finitely many real-analytic functions. A subset is subanalytic if it is locally the image, under a real-analytic map, of a relatively compact real-analytic subset (see [BierstoneMilman1988_Semianalytic-and-subanalytic-sets, Def. 3.1]). Many equivalent descriptions of subanalytic sets exist, see [BierstoneMilman1988_Semianalytic-and-subanalytic-sets, Prop. 3.13].
@@
The statement is local (by connectedness of the fibers) so by an appropriate choice of coordinates on $X$ we can assume that $\pi$ is a coordinate projection in a ball of $\mathbb{R}^{m+n}$ to $\mathbb{R}^{m}$. Using a multi-index notation with $(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, we can expand in Taylor series: $\rho(\mathbf{x},\mathbf{y})=\sum_{I}a_{I}(\mathbf{x})\mathbf{y}^{I}$ and then $Z$ is given as the vanishing locus of all the $a_{I}(\mathbf{x})$. However, by [Narasimhan1966_Introduction-to-the-theory-of-analytic-spaces, Ch. V, Cor. 2, pg. 100] it follows that $Z$ can be given as the vanishing locus of finitely many real-analytic functions in $\mathbf{x}$. ∎
@@ 3.2.4. Structure of real-analytic sets
By [BierstoneMilman1988_Semianalytic-and-subanalytic-sets, 2.11-12] an analytic set $Z$ has a stratification into finitely many locally closed pieces that are analytic submanifolds. The maximal dimension of a stratum is called the dimension of $Z$.
@@
Let $A^{\circ}\subset A$ denote the set of smooth points, i.e. it is real-analytic, with $\dim A^{\circ}=\dim A$, and $\dim\left(A\setminus A^{\circ}\right)<\dim A$ see e.g. [BierstoneMilman1988_Semianalytic-and-subanalytic-sets, Theorem 7.2]. Consider the map $\pi_{U}\colon I\to A$. Either $\dim A=\dim I$ (call this Case 1), or there exists a further relatively open subset $A^{\circ\circ}\subset A^{\circ}$ such that the map $\pi_{S}|_{I}$ has fibers of positive-dimension over $A^{\circ\circ}$ (call this Case 2).
@@
We have that $T_{a}UM=W^{s}(a)\oplus W^{c}(a)\oplus W^{u}(a)$ and $W^{c}(a)\subset T_{a}A$, since it represents the geodesic flow direction. In the quotient $T_{a}UM/W^{c}(a)$ the hyperplane $T_{a}A/W^{c}(a)$ must be transverse to at least one of the images of $W^{u/s}(a)$, say $W^{s}(a)$. Therefore the intersection $W^{cs}(a)\cap T_{a}A$ must have the expected dimension, so we have the equalities
@@
Pick an (n1)(n-1)-dimensional germ of smooth real-analytic submanifold $N\subset A^{\circ}$ with $a\in N$ such that $T_{a}UM=T_{a}N\oplus W^{cs}(a)$. For example, $N$ can be obtained by intersecting $A^{\circ}$ with a linear space of the appropriate dimension in local coordinates. We thus have that in a neighborhood of $a$, the manifold $N$ is transverse to the center-stable foliation.
@@
The map $\pi_{1}\colon\mathcal{W}^{cs}_{loc}[a]\to M$ is a diffeomorphism onto its image (see for example [BarreiraPesin2007_Nonuniform-hyperbolicity, §12.2]). Therefore, there exists an inverse map $\sigma_{a}\colon U\to\mathcal{W}^{cs}_{loc}[a]$ which is again a diffeomorphism onto its image. For $p\in U$, the unit tangent vector $\sigma_{a}(p)$ is characterized by the property that it is the unique vector above $p$ such that $g_{t}(\sigma_{a}(p))$ and $g_{t}a$ remain at bounded distance as $t\to+\infty$.
@@ 3.3. Analyzing the transitivity group
As we saw in § 2.2.14, there exists a continuous reduction of the structure group of the frame bundle to the transitivity group. Restricting this reduction to the unit sphere over one point of $M$ already yields nontrivial obstructions, first exploited by Brin and Gromov [BrinGromov1980_On-the-ergodicity-of-frame-flows]. A systematic analysis of known obstructions is given in [CekicLefeuvreMoroianu2023_On-the-ergodicity-of-the-frame-flow-on-even-dimensional-manifolds, §3.3], which we use to extract the needed constraints on the transitivity group. While ergodicity of the frame flow is not known in the generality we would need, the existing restrictions coupled with additional arguments are sufficient for our purposes.
@@ 3.3.1. The simplest obstruction
Recall that even-dimensional spheres have Euler characteristic $2$, hence do not admit a nowhere vanishing vector field, or even an everywhere defined tangent line-field. This is originally due to Poincaré in dimension $2$ [Poincare, Chapter 13] and Brouwer in higher dimensions [Brouwer]. In other words, if $n\geq 3$ is odd, there does not exist a continuous map
@@ 3.3.2. Surjectivity of image
We also record the following elementary consequence for even $n$. Suppose $\beta\colon\mathbb{S}^{n-1}\to\mathbb{S}^{n-1}$ is continuous, with the property that $\beta(e)\perp e$. Then $\beta$ is surjective when its image is descended to $\mathbb{R}\mathbb{P}^{n-1}:=\mathbb{S}^{n-1}/\pm 1$. Indeed, suppose by contradiction that the two vectors $\pm e_{n}$ are never in the image of $\beta$. Restricted to $\mathbb{S}^{n-2}(e_{n}^{\perp})$ we can decompose $\beta(e)=\beta_{1}(e)+\beta_{2}(e)$ with $\beta_{2}(e)=\alpha(e)e_{n}$ and $\beta_{1}(e)\in e_{n}^{\perp}$. Now by assumption $\beta_{1}(e)\neq 0$ and also $\left\langle e,\beta_{1}(e)\right\rangle=0$, so we can define
@@
which has the property that $\widetilde{\beta}(e)\perp e$. But this is a contradiction to § 3.3.1 since n2n-2 is even.
@@ 3.3.3. Restrictions on reduction of structure group
We next recall some more elaborate topological facts from [CekicLefeuvreMoroianu2023_On-the-ergodicity-of-the-frame-flow-on-even-dimensional-manifolds, Proof of Thm. 3.1, Step 1]. Consider the principal $\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$-bundle over the (n1)(n-1)-sphere:
@@ 3.3.4. Full frame bundle
We extend () 3.1.1 to include the full frame bundle:
[…]
@@
[…]
@@
With assumptions as in Theorem 1, let $H\subset\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$ be the transitivity group of $PM\to UM$, which we assume is connected by § 2.2.14. Then we have the following possibilities for $H$ and coset space $C:=\raisebox{-1.02228pt}{$H$}\backslash\raisebox{1.02228pt}{$\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$}/\raisebox{-1.02228pt}{$\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$}$:
@@
    $H=\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ and C=[1,1]C=[-1,1] (any $n$).
@@
    $H=\operatorname{\mathbf{SU}}(3)$ and C=[1,1]C=[-1,1] (for $n=8$ only).
Above, “any nn” means subject to $n\geq 3$ and the specified parity condition.
@@
Suppose next that $n$ is even. Because a hypersurface $S$ is necessarily odd-dimensional, we have that $H_{S}=\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ except possibly when $n=8$ in which case $H_{S}=\operatorname{\mathbf{SU}}(3)$ is also possible. So for all even $n\neq 8$ we have the claim, and since the orbits of $\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ on $\mathbb{S}^{n-2}\subset\mathbb{R}^{n-1}$ are classified by their last coordinate, an element of [1,1][-1,1]. For $n=8$ we also have the possibility that $H=H_{S}=\operatorname{\mathbf{SU}}(3)$. But the corresponding action of $\operatorname{\mathbf{SU}}(3)$ on $\mathbb{S}^{6}$ has orbit space again equal to [1,1][-1,1] and given by the value of the last coordinate. ∎
@@
From () 3.1.1 it suffices to show that the set $I(x)\subset FM(x)$ has the same dimension as $FM(x)$, since $I=\pi_{H}^{-1}(Z)$. Recall also that $FM=PM/\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$ and we have an action of $G_{t}$ lifting the geodesic flow $g_{t}$ on $UM$. Let $H\subset\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$ be the transitivity group of $PM\to UM$.
@@
    $\left.\raisebox{-1.00006pt}{$H$}\middle\backslash\raisebox{1.00006pt}{$\mathbb{S}^{n-2}$}\right.=\raisebox{-1.00006pt}{$H$}\backslash\raisebox{1.00006pt}{$\operatorname{\mathbf{SO}}_{n-1}(\mathbb{R})$}/\raisebox{-1.00006pt}{$\operatorname{\mathbf{SO}}_{n-2}(\mathbb{R})$}.$
@@
From § 3.3.4, only the cases when $C\neq\left\{pt\right\}$ need to be treated, i.e. C=[1,1]C=[-1,1] and $n$ is even.
@@
is a nonempty open set, as the intersection of two open spherical caps around $\beta(e_{1}^{\prime})$ and $\beta(e_{1}^{\prime\prime})$. Non-emptiness holds because the two spherical caps have measure adding up to the measure of $\mathbb{S}^{n-1}$, and the only situation in which they are disjoint is when $\beta(e_{1}^{\prime})=\beta(e_{1}^{\prime\prime})$. For a continuous path $e_{1}(t)$ with $e_{1}(0)=e_{1}^{\prime},e_{1}(1)=e_{1}^{\prime\prime}$, we have that any element $e_{n}\in L$ is in some $\mathbb{S}^{n-2}_{\alpha}(e_{1}(t))$, since $f(t):=\left\langle\beta(e_{1}(t),e_{n})\right\rangle$ satisfies $f(0)>\alpha,f(1)<\alpha$, so by the intermediate value theorem there exists $t$ such that $f(t)=\alpha$. This shows that an open set of hyperplanes belongs to $Z$ and concludes the proof. ∎
@@ 3.3.7. Last step
We know that $Z\subset HM$ is a closed real-analytic set of full dimension, therefore it is in fact all of $HM$. Indeed, for any point of $z\in Z$, the germ of $Z$ at $z$ denoted $Z_{z}$ has full dimension, therefore by [Narasimhan1966_Introduction-to-the-theory-of-analytic-spaces, Ch. V, Prop. 1, pg. 91] the complexified germ $\widetilde{Z}_{z}$ is an entire complex neighborhood of $z$, and hence the germ $Z_{z}$ is an entire real neighborhood of $z$ by the result just cited. It follows that $Z$ is also open, hence equal to $HM$.
It now follows from “Cartan’s axioms of $k$-planes” (with k=n1k=n-1) Theorem 2, that $M$ has constant sectional curvature. Since the geodesic flow is Anosov, the constant curvature must be negative.